OpenStudy (anonymous):

find the domain and range for the following equations: y=2x-1 y=(square root of -x)+8 y=x^2-6 y=2-1/x y=3(square root of x)+2 y=1/x+4 y=(square root of x+5) y=(x-6)^2 y=1/(x+4)^2

OpenStudy (anonymous):

that's a lot of equations

OpenStudy (anonymous):

Do you know how to find domain and range?

OpenStudy (anonymous):

sadly no. >.<

OpenStudy (anonymous):

Okay, take y = 2 - (1/x) for example

OpenStudy (anonymous):

what can x never equal?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

right? xD

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

so the domain is (-∞, 0) U (0,∞)

OpenStudy (anonymous):

always?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you so much. (:

OpenStudy (anonymous):

wait, no... just for that equation

OpenStudy (anonymous):

\[y=3\sqrt{x}\]

OpenStudy (anonymous):

what can x never equal?

OpenStudy (anonymous):

or, what does x always have to be greater than

OpenStudy (anonymous):

0

OpenStudy (anonymous):

can it be 0?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes! so, the domain is [0, ∞)

OpenStudy (anonymous):

i dont remember the difference between the ( and [

OpenStudy (anonymous):

[ means it includes ( it doesn't include since we can never reach ∞, ( will always be used

OpenStudy (anonymous):

same with -∞

OpenStudy (anonymous):

the range is a bit complicated.

OpenStudy (anonymous):

oh god ><

OpenStudy (anonymous):

lets go back to: y = 2 - (1/x)

OpenStudy (anonymous):

x can never be 0, so if we look on the graph: http://www4a.wolframalpha.com/Calculate/MSP/MSP20311hg12eagf2a7hh490000541c8369027g5cc5?MSPStoreType=image/gif&s=62&w=325.&h=151.&cdf=RangeControl what value can y never be?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Hint: It's 2

OpenStudy (anonymous):

the y value

OpenStudy (anonymous):

WHAT this confuses me xD

OpenStudy (anonymous):

set y = 2 and solve this for x: y = 2 - (1/x)

OpenStudy (anonymous):

that's impossible

OpenStudy (anonymous):

so the range will be (-∞, 2) U (2, ∞)

OpenStudy (anonymous):

im still confused but thank you for the help. (:

OpenStudy (anonymous):

\[y=\sqrt{x+5}\]

OpenStudy (anonymous):

what must x be greater than?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

in this equation x must be greater than or equal to -5 because you can't have negative square roots

OpenStudy (anonymous):

so the domain will be [-5, ∞)

OpenStudy (anonymous):

omg i actually get it :o

OpenStudy (anonymous):

so, the range will be [0, ∞)

OpenStudy (anonymous):

i just dont get how thats the range.